# Nonlinear chiral metasurfaces based on structured van der Waals materials

### v1

@@ Abstract
Nonlinear chiral photonics explores nonlinear response of chiral structures, and it offers a pathway to novel optical functionalities not accessible through linear or achiral systems. Here we present the first application of nanostructured van der Waals materials to nonlinear chiral photonics. We demonstrate the three orders of magnitude enhancement of the third-harmonic generation from hBN metasurfaces driven by quasi-bound states in the continuum, accompanied by a strong nonlinear circular dichroism in the resonant metasurface varying from the values -0.5 to +0.5. This novel platform for chiral metaphotonics can be employed for achieving large circular dichroism combined with high-efficiency harmonic generation in a broad frequency range.
@@ I Introduction
In this paper, we study dielectric metasurfaces to enhance nonlinear response and present the first application of hBN-based metasurfaces in nonlinear chiral photonics, demonstrating the generation of chiral optical harmonics through nonlinear optical processes. We engineer and fabricate nonlinear hBN metasurfaces supporting chiral quasi-BIC mode in the near-IR range. We predict theoretically and demonstrate experimentally three orders of magnitude enhancement of the third harmonic generation compared to the thin film driven by quasi-BIC. We show that nonlinear CD changes from the value +0.5 to -0.5 with pump wavelength. We believe the exploration of hBN in this capacity underscores its potential to contribute significantly to the development of advanced photonic devices and systems and our results pave the way toward a new platform for biosensing in the whole visible range and near UV based on nanostructured van der Waals materials.
@@ II Results and Discussion
Our metasurfaces consist of meta-atoms placed on top of fused silica substrate arranged into a square lattice, as shown in Figure 1a. The unit cell is two bars, one of which is shorter in length and rotated on an angle relative to the axis between the bars. We engineer the metasurface to support quasi-BIC mode in the near-infrared spectral range. The geometrical parameters of our meta-atoms are nm, nm, nm, nm, , nm, nm, as shown in Figure 1b. For the non-perturbed case, i.e. for and , the resonant mode of interest is going to have an infinite -factor. We fabricate the metasurface from hBN film, and the scanning electron microscopy (SEM) image of the metasurface obtained is shown in Figure 1c.
@@
The experimentally obtained transmission spectra for co-polarized and cross-polarized light are shown in Figure 2d (for more details of the experiment see Methods and Supporting Information S1). The spectra for RCP and LCP co-polarised light demonstrate the dip at the wavelength of 1494 nm and no significant difference between them. Meanwhile, the cross-polarized transmission spectra show a value different from the zero only in the vicinity of the resonance. The behaviour observed is in good agreement with theoretical prediction.
@@
Next, we simulate and measure the THG for the designed hBN metasurface and evaluate the nonlinear CD. For nonlinear simulation, we use the undepleted pump approximation. The THG signal is calculated in the transmission direction in the zero diffraction order. The spectra obtained for co-polarised pump light and harmonic signal with RCP (RR) and LCP (LL) are shown in Figure 3a by blue and red curves, respectively. We also acquired the full THG signal generated from the structure pumped by LCP and RCP, as shown in Figure 3a by purple and yellow curves, respectively. The simulated THG signal exhibits enhancement of six orders of magnitude at the pump wavelength of 1494 nm compared to the signal from the pump outside the resonance. The zoomed-in area near the resonance is shown in Figure 3a inset. One can see that the amplitude of the THG signal depends on the pump polarization. To evaluate the nonlinear CD we use the two terms taking into account only co-polarised signal () and the full signal pumped () defined as
@@
where is the intensity of the forward generated LCP(RCP) THG signal with LCP(RCP) pump polarization, is the full forward intensity of THG signal with LCP(RCP) pump polarization. The resulting nonlinear CD curves are shown in Figure 3b. Both curves exhibit a similar trend. The maximum of nonlinear CD value is observed at the wavelength of 1475 nm. However, this value corresponds to a low THG signal outside the resonance, and at the vicinity of the resonance, the nonlinear CD value changes from 0 to +0.35.
@@
We further measure THG from fabricated hBN metasurfaces pumped in the near-IR range from 1400 nm to 1600 nm (for more details see Methods and Supporting Information S1). The spectrum of the THG pumped at a wavelength of 1490 nm has a maximum at a wavelength of 497 nm and is shown in Figure 4a. The power dependence recorded has a good agreement with cubic as shown in the inset of Figure 4a. The maximum THG efficiency of 10-6 is observed at the pump wavelength of 1490 nm and the peak intensity of 10 GW/cm2. We notice that only zero diffraction order of THG is collected, so the full THG efficiency tends to be higher.
@@
To investigate the nonlinear CD, we pump the metasurfaces with LCP and RCP and separate the LCP and RCP of THG in the collection. We pump the metasurface with wavelengths from 1400 nm to 1600 nm and record the THG spectra. Next, we determine the maximum of the THG for every pump wavelength. The maximum of the THG for different pump wavelengths is shown in Figure 4b. The THG signal strongly depends on the pump wavelength. At the resonance, the THG signal is enhanced by about three orders of magnitude compared to the THG signal from the thin hBN film (see Supporting Information S2). Based on Eqs (2) and (3), we calculate nonlinear CD for only co-polarised components (CD(3ω)) and total nonlinear CD without separation of LCP and RCP in the collection (CD). The resulting dependencies are shown in Figure 4c. The total nonlinear CD without separation of LCP and RCP in the collection shows the switching of the value from -0.4 to +0.2 near the resonance of the structure, whereas the nonlinear CD of the only co-polarised components exhibits the change from -0.5 to +0.5 near the resonance.
@@
In addition, we observe the second-harmonic generation (SHG) from the metasurfaces in the vicinity of the resonance. The SHG is several orders of magnitude lower than the THG and the power dependence demonstrates a good agreement with quadratic law. The SHG pumped at the resonance wavelength exhibits different amplitude for LCP and RCP pumps. To estimate the nonlinear CD for SHG, we used the similar definitions, and , where is the intensity of the forward generated LCP(RCP) SHG signal with LCP(RCP) pump polarization, is the full forward intensity of THG signal with LCP(RCP) pump polarization. Both nonlinear CDs for SHG acquired according to the equations show a positive value. The total CD at the resonance, whereas for co-polarised collection CD (for more details, see Supporting Information S3).
@@ III Conclusions
In conclusion, we have observed the three orders of magnitude enhancement of the third harmonic generation from hBN metasurfaces driven by quasi-BIC mode. We have experimentally demonstrated the strong chiral response of the metasurface changes from -0.5 to +0.5. We have introduced a new platform for chiral photonics, based on nanostructured van der Waals materials supporting optical resonances, which can be employed for chiral biosensing in the whole visible range and near UV.

### v2

@@ Abstract
Nonlinear chiral photonics explores nonlinear response of chiral structures, and it offers a pathway to novel optical functionalities not accessible through linear or achiral systems. Here we present the first application of nanostructured van der Waals materials to nonlinear chiral photonics. We demonstrate the three orders of magnitude enhancement of the third-harmonic generation from hBN metasurfaces driven by quasi-bound states in the continuum and accompanied by strong nonlinear circular dichroism at the resonances. This novel platform for chiral metaphotonics can be employed for achieving large circular dichroism combined with high-efficiency harmonic generation in a broad frequency range.
@@ I Introduction
In this paper, we study dielectric metasurfaces to enhance nonlinear response and present the first application of hBN-based metasurfaces in nonlinear chiral photonics, demonstrating the generation of chiral optical harmonics through nonlinear optical processes. We engineer and fabricate nonlinear hBN metasurfaces supporting quasi-BIC mode in the near-IR range. We predict theoretically and demonstrate experimentally three orders of magnitude enhancement of the third harmonic generation compared to the thin film driven by quasi-BIC. We show that nonlinear CD changes from the value +0.12 to -0.38 with pump wavelength. We believe the exploration of hBN in this capacity underscores its potential to contribute significantly to the development of advanced photonic devices and systems and our results pave the way toward a new platform for biosensing in the whole visible range and near UV based on nanostructured van der Waals materials.
@@ II Results and Discussion
Our metasurfaces consist of meta-atoms placed on top of fused silica substrate arranged into a square lattice, as shown in Figure 1a. The unit cell is two bars, one of which is shorter in length and rotated on an angle relative to the axis between the bars. The geometrical parameters of our meta-atoms are nm, nm, nm, nm, , nm, nm, as shown in Figure 1b. We engineer the metasurface to support quasi-BIC mode in the near-infrared spectral range. We fabricate the metasurface from hBN film, and the scanning electron microscopy (SEM) image of the metasurface obtained is shown in Figure 1c (for fabrication process details see Supporting Information S1).
@@
The experimentally obtained transmission spectra for co-polarized and cross-polarized light are shown in Figure 2d (for more details of the experiment see Supporting Information S2). The spectra for RCP and LCP co-polarised light demonstrate the dip at the wavelength of 1494 nm and no significant difference between them. Meanwhile, the cross-polarized transmission spectra show a value different from the zero only in the vicinity of the resonance. The behaviour observed is in good agreement with theoretical prediction.
@@
Next, we simulate and measure the THG for the designed hBN metasurface and evaluate the nonlinear CD. For nonlinear simulation, we use the undepleted pump approximation. The THG signal is calculated in the transmission direction in the zero diffraction order (for more details see Supporting Information S3). The spectra obtained for co-polarised pump light and harmonic signal with RCP (RR) and LCP (LL) are shown in Figure 3a by blue and red curves, respectively. We also acquired the RCP THG signal generated from the structure pumped by LCP (RL) and the LCP THG signal generated from the structure pumped by RCP (LR), as shown in Figure 3a by purple and yellow curves, respectively. The simulated THG signal exhibits enhancement of six orders of magnitude at the pump wavelength of 1494 nm compared to the signal from the pump outside the resonance. The zoomed-in area near the resonance is shown in Figure 3a inset. One can see that the amplitude of the THG signal depends on the pump polarization. To evaluate the nonlinear CD we use the two terms taking into account only co-polarised signal () and the full signal pumped () defined as
@@
where is the intensity of the forward generated LCP(RCP) THG signal with LCP(RCP) pump polarization, is the full forward intensity of THG signal with LCP(RCP) pump polarization. The resulting nonlinear CD curves are shown in Figure 3b. Both curves exhibit a similar trend. The maximum of nonlinear CD value is observed at the wavelength of 1475 nm. However, this value corresponds to a low THG signal outside the resonance, and at the vicinity of the resonance, the nonlinear CD value changes from -0.35 to 0.
@@
We further measure THG from fabricated hBN metasurfaces pumped in the near-IR range from 1400 nm to 1600 nm (for more details see Supporting Information S2). The spectrum of the THG pumped at a wavelength of 1490 nm has a maximum at a wavelength of 497 nm and is shown in Figure 4a. The power dependence recorded has a good agreement with cubic as shown in the inset of Figure 4a. The maximum THG efficiency of 10-6 is observed at the pump wavelength of 1490 nm and the peak intensity of 10 GW/cm2. We notice that only zero diffraction order of THG is collected, so the full THG efficiency tends to be higher.
@@
To investigate the nonlinear CD, we pump the metasurfaces with LCP and RCP and separate the LCP and RCP of THG in the collection. We pump the metasurface with wavelengths from 1400 nm to 1600 nm and record the THG spectra. Next, we determine the maximum of the THG for every pump wavelength. The maximum of the THG for different pump wavelengths is shown in Figure 4b. The THG signal strongly depends on the pump wavelength. At the resonance, the THG signal is enhanced by about three orders of magnitude compared to the THG signal from the thin hBN film (see Supporting Information S4). Based on Eqs 2 and 3, we calculate nonlinear CD for only co-polarised components (CD(3ω)) and total nonlinear CD without separation of LCP and RCP in the collection (CD). The resulting dependencies are shown in Figure 4c. The total nonlinear CD without separation of LCP and RCP in the collection shows the switching of the value from -0.38 to +0.12 near the resonance of the structure, whereas the nonlinear CD of the only co-polarised components exhibits the change from -0.35 to +0.12 near the resonance. The difference between the numerically simulated spectra and the experimental results can be attributed to at least two factors: (1) side walls of fabricated bars are not exactly vertical but slightly inclined (see Fig. 1c); and (2) in numerical simulations we assumed that the crystallographic axes are aligned with the metasurface grating direction, which could be different in the real structure, and extremely hard to control.
@@
In addition, we observe the second-harmonic generation (SHG) from the metasurfaces in the vicinity of the resonance. The SHG is several orders of magnitude lower than the THG and the power dependence demonstrates a good agreement with quadratic law. The SHG pumped at the resonance wavelength exhibits different amplitude for LCP and RCP pumps. To estimate the nonlinear CD for SHG, we used the similar definitions, and , where is the intensity of the forward generated LCP(RCP) SHG signal with LCP(RCP) pump polarization, is the full forward intensity of THG signal with LCP(RCP) pump polarization. The total second harmonic CD shows the value of -0.27 at the resonance of the structure, whereas for co-polarised collection CD (for more details, see Supporting Information S5).
@@ III Conclusions
In conclusion, we have observed the three orders of magnitude enhancement of the third harmonic generation from hBN metasurfaces driven by quasi-BIC mode. We have experimentally demonstrated strong nonlinear chiral response of the metasurface at the resonances. We believe our results introduce a new platform for chiral photonics, based on nanostructured van der Waals materials supporting optical resonances, which can be employed for chiral biosensing in the whole visible and near UV frequency ranges.